• 8 August 2019: SINFONI has been decommissioned at the end of June 2019. The SINFONI entry has been removed from Phase 1.
  • 21 April 2019: The SINFONI intervention was successful and the MACAO is repaired with a new membrane. SINFONI is now fully operational.
  • 10 April 2019: Cooling water was detected inside MACAO that came from a leak in the cooling lines of the Cassegrain TCCD. Because of this leaking problem, SINFONI has been dismounted from UT3. Service and visitor observations are therefore suspended until MACAO is fixed.
  • 13 March 2019: SINFONI will NOT be offered in Period 104.
  • 26 Febuary 2019: The SINFONI transfer was successful, and the recommissioning was carried out in the first two weeks of February. After moving to UT3, the system performance is very similar to that at UT4. Some changes include: 1). the slit positions are shifted by 4 pixels in the positive x-direction; 2). the detector blind spot is  shifted by 4 pixels more into slit #16, which is similar to the previous position before the 2018 intervention.
  • 20 January 2019: SINFONI will be removed from UT4 and transfered to UT3. Recommisioning of SINFONI is scheduled at the end of January. SINFONI is expected to be available for science operation in early Febuary. A summary of the SINFONI transfer will be available after SINFONI is operational at UT3.
  • 19 Febuary 2018: The SINFONI intervention was successful, and the recommissioning was carried out in the beginning of February. The damaged preoptics motor has been replaced and the new motor is functioning well. Maintenance tasks were also performed on many parts of SINFON, such as the MACAO AO system. After the intervention, the system performance is, overall, similar to that before the intervention. Some changes include:        
  1. Installation of a new membrane mirror. Slight improvement in internal focus in shorter wavelengths (J and H) with no impact in K.
  2. The orientation of SINFONI with respect to the sky has been recalculated and numerically adjusted by 0.26 degrees for the NGS mode and 0.5 degrees for the LGS mode.
  3. The locations of the spectra are shifted slightly on the detector, by about 4 pixels in x and 1-2 pixels in y.
  4. The detector blind spot is now shifted 4 pixels farther away from the center of the field of view, while shifting more into slit #1 (bottom slit in the FOV).
  5. The pipleline reduces the new data without issues, except for the distortion recipe for the J-band NorthSouth data, where the bp_dist-min_cut parameter should be set to 0.05 (instead of 0.1). This value is found to work for data before and after the intervention, and will be implemented in the next release of the pipeline.
  • 5 January 2018: SINFONI intervention has been moved forward. SINFONI will be offline from 16 Jaunary onwards until about 7 February.
  • 30 November 2017: SINFONI is foreseen to undergo an intervention to replace the preoptics wheel motor, and other maintenance tasks, in mid-February.
  • 30 November 2017: New version of User Manual (v101.1) is available.
  • 04 November 2017: Recently UT4 has been experiencing intermittent issues with image elongation, affecting the data taken by SINFONI. While the cause of this occasional elongation is under investigation, the following rules will be applied for SINFONI OB classification:
    1) For AO observations (NGS, LGS), SINFONI data is classified by the Strehl ratio for AO, no ellipticity criterion is applied.
    2) For noAO observations and the object is a stellar object and FWHM can be reliably measured, then:
            If elongation < 10% for most stars then class as A
            If 10% < elongation < 20% for most stars class as B
            If 20% < elongation class as C
    If no compact objects are in the field-of-view, then a telluric star taken before or after will be used to estimate elongation. The criteria same as above, except the numbers are relaxed to 15% and 25% instead of 10% and 20%.
  • 14 September 2017: New version of User Manual (v101.0) is available.
  • 30 June 2017: The laser is now is now back in operation with SINFONI.
  • 18 June 2017: Due to a technical problem, SINFONI observations with the laser will not be available until at least the end of June.
  • 7 March 2017: New version of the User Manual (v100.1) is available.
  • 4 December 2016: SINFONI underwent successfully recommissioning following the installation of the deformable secondary mirror on UT4.
  • 23 September 2016: SINFONI will be offline from approx 9 October onwards until the end of November due to M2 installation.
  • 13 September 2016: The MACAO shutter for the AO unit was replaced successfully in a mini intervention on 12/9. 
  • 14 March 2016: As SINFONI has recently undergone upgrade and the manual is being updated with new information, for the Call for Proposals please use the Period 97.0 manual, in conjunction with these supplementary notes. In a nutshell, the performance improvements include:
    •  A significant improvement in the overall J-band throughput
    • Increased spectral resolution in all bands
    • A spot of dead pixels which previously affected the centre of the field of view are now shifted mostly out of the way
    • The vignetting of slit 32 at the top of the field of view is now much reduced
    • An improved AO performance of ~10% in Strehl
    • No significant negative impact on the system performance.
  • 1 December 2015: SINFONI upgrade will take place during January 2016. It will be dismounted from UT4 on 20 December. Recommissioning will take place at the end of January, afterwhich the instrument will be back in normal service.
  • 11 August 2015: SINFONI will not be available during the following periods due to technical activities:
    • August 26th to September 1st: re-commissioning of HAWKI after the GRAAL installation, and test with the 1st unit of 4LGSF.
    • September 21 to September 30th: UT4 maintenance
  • 29 May 2015: New version of the User Manual (v96.1) is available. To be used for Phase 2 preparation.
  • 26 February 2015: The P96 Call for Proposals is out. The ETC has also been updated. Note the following:
    • An upgrade of the SPIFFI optics is foreseen to take place during Period 96. Its scope and schedule are being finalized as this call is released. Improvements may include replacing J and K filters by new ones with higher transmission, replacement of the pre-optics by one with better transmission in the J band, and replacement of mirrors by ones with better surface quality to improve the line spread functions. The instrument will therefore not be available for approximately 1 month likely to start in January 2016.
    • The installation of the Deformable Secondary Mirror and re-commissioning of the telescope is expected to take place in October 2016; as a consequence, UT4 will not be available for approximately 2 months.
    • The commissioning of the AOF and AOF-related systems is expected to require monthly slots of up to 7 days during possibly 2 years.
  • 27 August 2014: The P95 Call for Proposals is out.
  • 2 June 2014: grating wheel speeds have been lowered: the normal speed is 14 encoders/ms, and the fine speed is 5 enc/ms. A grating change between the 2 most separated gratings (J and H+K) will now take 3.7 minutes.
  • 13 April 2014: grating wheel is back to normal functionality, albeit with reduced speed.
  • 3 April 2014: SINFONI is experiencing problems with the grating wheel. As a result only one grating position is allowed per day, set during daytime.
  • 5 March 2014: SINFONI may undergo an upgrade during Period 96 -- at the same time as the installation of the Deformable Secondary Mirror -- to replace the  H+K grating by a high-resolution one, to replace the detector, and to modify its fore-optics, in prevision of its later insertion into ERIS.
  • 27 February 2014: New version of manual (v94.0) uploaded. To be used for period 94.
  • 9 January 2013: New version of manual (v93.2) uploaded. To be used for Phase 2 preparation.
  • 2 October, 2013: New version of manual (v93.1) uploaded.
  • 28 September, 2013: Additional information on the new pupil tracking mode is available in this document. (Note added later: this is superceeded by information in subsequent versions of the User Manual).
  • 21 September, 2013: SINFONI recommissioning was successful. All filter/grating combinations and observing modes are again available. Calibration database is being repopulated. Dark exposures and linearity tests are resumed.
  • 17 September, 2013: SINFONI intervention was successful. Four motors were replaced (grating, filter, pre-optics and sky-spider). The coating of the 25mas pre-optics lens was discovered to be degraded (likely during warm up) and this was successfully polished off. The detector is displaced 5 pixels compared to before (slit positions are shifted 5 pixels in the positive x-direction.)
  • 2 September, 2013: A new Pupil-Tracking mode will be offered in Period 93, enabling the Spectrocopic Angular Diffrential Imaging (S-ADI) technique to be used for certain targets. Please see the SINFONI Manual for more details.
  • 1 September, 2013: SINFONI intervention underway. Instrument will be back online in the second half of September.
  • 24 July, 2013: Filter wheel is still stuck in K. Due to the delicate state of the filter wheel and the substantial risk of getting it stuck while moving it (thus rendering the instrument inoperable), the Instrument Operation Team has decided to keep SINFONI in K until 12 August, afterwhich we will attempt to move it to H+K. Intervention is scheduled in the beginning of September during the M1 recoating. We apologise for any invonvenience caused.
  • 10 June, 2013: Filter wheel is still stuck in K, and the aforementioned shift in wavelength is still happening.
  • 12 May, 2013: The grating has displaced slightly, resulting in a shift of 2-3 pixels in the dispersion direction compared to normal. Since the science and the arcs are taken with the same shift, the data can be calibrated.
  • 07 May, 2013: Filter wheel is stuck in K; the situation is being assessed.
  • 26 April, 2013: Filter wheel movement has been restored. Service mode observations are resumed.
  • 15 April, 2013: We are experiencing problems with filter wheel. Currently the only solution is to keep the filter in a fixed position under simulation mode. As a result darks exposures cannot be taken. Service observations are therefore suspended until this problem is fixed.
  • 2 April, 2013: LGS has a shutter problem and is not available until the problem is fixed.
  • 1 March, 2013: The new laser PARLA has been successfully commissioned and is now in operation with SINFONI, with LGS available on most nights. The on-sky power is about 6W.
  • 26 Feb, 2013: The strong noise pattern as reported in November 2012 is still occassionally seen. The pattern is not fixed in position, but moves around the detector, with a RMS > 10 ADUs. Observers should look out for this.
  • 20 Feb, 2013: The motors for the Grating and Filter wheels are failing. Intervention is expected in September 2013. In the meantime, in order to prolong the lives of the motors and minimise further damage, the speed of the motors have been drastically decreased, such that a grating/filter change will now take about 10 minutes. While it is still possible to change grating/filter during the night, the user is enouraged to stick to one grating/filter setting per night. Observers should adapt their observing strategies in light of this.
  • 29 November, 2012: Following troubleshooting by Instrumentation, the strong noise pattern has mainly gone. Service mode OBs with faint targets are put back into the queue.
  • 11 November, 2012: Noise pattern is seen in science and calibration data. This is being investigated. Until this problem is solved, for service mode observations preference will be given to bright targets.
  • September, 2012: The manual for period 91 is unchanged with respect to P89 and P90. An update of the manual will be expected by October 2012.
  • March 21 to May 3, 2012: Because of preparatory work for the installation of the Adaptive Optics Facility, Yepun is not available for operations. Observations with SINFONI are therefore not possible.
  • February 2012: The user manual is unchanged for P90 (with respect to P89).
  • August 2011: The user manual was updated for P89; in particular, details on detector persistence were added (please see section 2.6).
  • December 2010: With the start of P87 Phase 2, the new template 'SINFONI_ifs_cal_GenericOffset' has been made available, to allow efficient PSF observations
    in LGS-noTTS mode (Seeing Enhancer mode); via this template, PSF calibrations data can be taken within the same OB as the science data.
    Use of this template is restricted to LGS-noTTS mode observations.
    For more details please see the updated P87 version of the user manual, in particular section 12 (template description).
  • August 2010: A SPIE article (Mirny et al., 2010, SPIE Proceedings, Vol. 7737, p. 77371R) has been published to describe the improvements recently implemented in the SINFONI data reduction pipeline as already mentioned below (please see February 2010 news).
  • May 2010: For LGS mode operations, SINFONI has received a new STRAP head that should help to overcome the problem with the central obscuration of the old STRAP head experienced under very good seeing conditions (i.e., due to its narrow PSF, the TTS was lost in the central obscuration).
    Work is on-going to characterize the performance improvements of the new STRAP head; SINFONI LGS mode users, please check the instrument news as information will be provided as soon as it is available.
  • February 2010: An updated version of the SINFONI data reduction pipeline was released, featuring two significant improvements: (i) the implementation of a distortion and dispersion correction, and (ii) the implementation of a possibility to monitor instrument efficiency and stellar zero-points.
  • January 20, 2010: The user manual has seen some minor updates.
  • June 2008: Please refer to the Observatory News page for the latest update on LGS (un-)availability.
  • September 2007: The user manual has been updated to include: (i) new R-band magnitude limit of R=18 for the TTS brightness in LGS mode, and (ii) an update on the orientation of STRAP field selector.
  • April 2007: The official VLT Laser Guide Star Facility Policies have been released by Paranal Observatory Science Operations.
  • March 2006: The user manual for period 78 has been updated to reflect the availability of the Laser Guide Star and Rapid Response Modes.
  • January 2006: The user manual for period 77 is unchanged with respect to period 76. The page with frequently asked questions was updated with more hints to query the 2MASS and other catalogs from skycat. A direct link to the SINFONI data reduction pipeline as released in October 2005 was add to the tools.
  • July 2005: More informations for visiting astonomers about SINFONI data are available.
  • June 2005: Updated user manual for period 76 service and visitor mode observation runs and the P77 call for proposals. This version includes updated performance estimates of SPIFFI and SINFONI-AO.
  • April 2005: The page with frequently asked questions was updated with informations on how stacks of image cubes of very faint targets can be coadded accuractely.
  • April 2005: The data taken with SINFONI are now visible from the ESO Archive.
  • ... older messages